# EVEREST [Uncaptioned image]: Efficient Masked Video Autoencoder by Removing Redundant Spatiotemporal Tokens

### v4

@@ 1 Introduction
Massive video data floods the web and media daily with the rapid growth of portable devices equipped with cameras, such as AR glasses, smartphones, UAVs, and robots. However, direct utilization of user-generated video data for solving target task problems is nontrivial, as annotating them is time-consuming and expensive. One potential approach to tackle this problem is to learn generic representations from unlabeled video streams that can transfer to diverse downstream visual tasks. Video Representation Learning (VRL) (Fernando et al., 2017; Piergiovanni et al., 2020; Qian et al., 2021; Pan et al., 2021) methods allow learning spatial and temporal features from input video frames in a self-supervised manner without any human annotations. A caveat to such pre-training for video tasks is that, unlike image-based tasks with static information of objects in instances, video-based tasks involve temporal causality; that is, successive frames are strongly correlated in their semantics.
@@
Recently, Masked Video Autoencoder (MVA) (Tong et al., 2022; Feichtenhofer et al., 2022), which learns to reconstruct randomly masked spatiotemporal regions in video clips, has shown impressive performance on various video-based problems. Yet, critical challenges remain in efficiently exploiting the spatiotemporal information in real-world videos: (1) Tokens (a pair of two temporally successive patches in the same space) in videos are not equally valuable to reconstruct, as the amount of their information not only depends on spatial importance but also on temporal redundancy. (2) Learning representations from videos is infeasible without a huge computing budget. MVA approaches (Feichtenhofer et al., 2022; Wang et al., 2023b, c) that reconstruct the whole video require excessively large amounts of computations, making it impractical without access to a substantial GPU cluster. For example, VideoMAE (Tong et al., 2022) takes about 27 hours to pre-train for 800 epochs with ViT-B using 64 NVIDIA V100 (32GB) GPUs, and ST-MAE (Feichtenhofer et al., 2022) takes about 35.8 hours to pre-train for 800 epochs with ViT-L using 128 NVIDIA A100 (80GB) GPUs.
Refer to caption (a) Computation (GFLOPs) Comparison
Refer to caption (b) Memory comparison
Figure 1: Efficiency of our EVEREST against VideoMAE on K400 dataset. (a) GFLOPs for pre-training and fine-tuning. The bubble size is proportional to the GFLOPs of the model. (b) Memory consumption using one node equipped with 8$\times$ A100 (80GB). VideoMAE with ViT-B and -L fails to deploy the model due to out-of-memory if the batch sizes are 512 or larger. For a ViT-L backbone with a batch size of 256, our method achieves about 4×4$\times$4 × less memory consumption than VideoMAE. Please see  Tabs. 1 and 3 for detailed results.
@@
To overcome such limited feasibility of video representation learning, we propose an Efficient Masked Video AutoEncoder which removes REdundant Spatiotemporal Tokens (EVEREST), that is a highly efficient training algorithm with token selection for alleviating the extremely high resource requirements for VRL. Unlike prior methods (Tong et al., 2022; Feichtenhofer et al., 2022; Wang et al., 2023a) reconstructing all patches, we aim to select a subset of informative visual tokens to learn based on the distance disparity across temporally adjacent tokens in the embedding space. Our redundancy-robust token selection approach successfully detects meaningful changes in the state/object of incoming videos in an online manner, discarding less meaningful tokens containing redundant information or meaningless backgrounds, without resorting to dense motion features from the incoming video, such as HOG or optical flows (Sun et al., 2023; Wang et al., 2023c).
@@
This allows us to significantly reduce computational cost and memory usage while maintaining the quality of a representation model by back-propagating to only a few selected tokens that retain rich spatiotemporal information with minimal redundancy. As shown in Fig. 1, our EVEREST saves computation costs by $\bf 26\sim 45$in pre-training and $\bf 44\sim 48$in fine-tuning across varying ViT scales while achieving competitive performance against strong VideoMAE baselines. We also achieve an impressive amount of memory reduction with all ViT backbones, which becomes more effective with larger models. For example, on a ViT-Large with a batch size of 256, our method achieves about $\bf 4\times$ smaller memory consumption than VideoMAE and enables single-node training with larger batch sizes using a large backbone, whereas existing VRL methods require immense memory occupancy and fail to train in the same setup as they go out-of-memory.
@@ 4.1 Redundancy-robust Mask Generation
$arg\min_{\mathcal{W},\bm{w}}~{}\!\sum^{N}_{n=1}\sum^{\tau-1}_{i=0}\left\|f_{\mathcal{W}}({(\widetilde{\bm{k}}_{i}^{\prime})^{n}})-{{\bm{m}_{i}^{n}\otimes\bm{v}_{n}}}[2i:2i+1]\right\|_{F},$ (3)
$\text{where}~{}~{}\bm{m}^{n}\!=\!G\left([J\cdot\rho_{pre}],\rho_{post},\tau\right)\in\{0,1\}^{\tau\times[J\cdot\rho_{pre}]}.$
We omit the positional embedding term for notational simplicity. Whereas earlier MVA works used a fixed masking ratio per video cues for self-supervised training, our approach allows a dynamic masking rate for each frame by design, based on the occupancy of essential tokens. As shown in Fig. 6, each video frame contains a different amount of spatiotemporal information, and our dynamic masking strategy enables the model to focus on learning the valuable representation in a holistic view from incoming videos. We note that our EVEREST can be generalized well on both pre-training and fine-tuning phases with a negligible additional computational cost and training time (Please see Fig. 1). We find that removing unimportant tokens is crucial for both phases, obtaining more meaningful representations and achieving superior performances. (Please see Tab. 10). The overall process of our redundancy-robust token selection is illustrated in Fig. 2.
@@
Figure 4: Illustration of information-intensive frame selection. We adaptively select frames based on the ReRo token frequency, which indicates significance compared to frames.
Table 1: Performance Comparison on K400. PT and FT mean pre-training and fine-tuning, respectively. We use an input size of $16\times 224^{2}$. Memory usage is measured with 256 batch size when pre-training. ${}^{\dagger}$ is trained from random initialization. VideoMAE and EVEREST are pre-trained for 200 epochs, and we refer (Wightman, 2019) for computing ${}^{\ddagger}$.
@@
MViT (Fan et al., ${}^{\dagger}$ MViT-S - 32.9 - 76.0
MViT (Fan et al., ${}^{\dagger}$ MViT-B - 70.5 - 78.4
ViViT FE (Arnab et al., 2021) ViT-L IN-21K ${}^{\ddagger}$ 3980.0 N/A 81.7
K-centered (Park et al., 2022) XViT IN-1K ${}^{\ddagger}$ 425.0 N/A 73.1
K-centered (Park et al., 2022) Mformer IN-1K ${}^{\ddagger}$ 369.5 N/A 74.9
K-centered (Park et al., 2022) TSformer IN-1K ${}^{\ddagger}$ 590.0 N/A 78.0
@@ 4.2 On-the-fly Information-intensive Frame Selection
Table 3: Memory usage comparison during the pre-training (one node, A100$\times$8 GPUs) on K400. The memory gaps grow up when increasing batch size and model size. Especially, our EVEREST with ViT-L achieves about $\mathbf{4\times}$ better efficiency than VideoMAE (634.1GB). OOM indicates out-of-memory.
@@ 5 Experiments
Table 5: EVEREST-Finetuning with other MVAs on K400. We apply our EVEREST for finetuning the pre-trained models (ViT-B) by SoTA MVAs. We measure the memory usage with the same batch size of 128. While SoTA methods use full tokens during fine-tuning, our EVEREST uses only redundancy-robust tokens that reduces significant computational cost and memory usage with comparable accuracy.
@@
EVEREST has a potential for estimating motion information. We provide masked input examples with their importance heatmaps through our proposed redundancy-robust masking in Fig. 6. Our masking strategy enables the model to capture the most informative tokens containing static and dynamic moving of objects (e.g., doing wall push up (Fig. 6 (a)) and playing ping pong (Fig. 6 (b)). Furthermore, our masking strategy captures informative objects even with the rapid change of the view in the first-person videos, as it masks out objects which are crucial for understanding the camera wearers’ attention (Fig. 6 (c)), while not attending to backgrounds such as walls and floors.
@@
EVEREST adaptively finds information-dense frames from egocentric videos. Real-world videos include many redundant scenes over a long time, the VRL model may waste time and computations by learning on meaningless frames, which also can lead to poor local optima due to bias and catastrophic forgetting. To overcome the limitation, we adopt the rate of redundancy-robust sampling $\alpha$ to focus on the frames with larger motions across frames as illustrated in Fig. 3. By constructing the given videos to have fluent motion information, our model focuses more on learning the core part of the video. We set the default $\alpha$ to 1.5 and report the effect of $\alpha$ in Appendix C.We quantitatively compare our method to recent SoTA methods using OSCC task on Ego4D in Tab. 6. We pre-train OSCC without labels and fine-tune it to classify whether the object’s state has changed. With only visual information, we outperform the previous SoTA method Egocentric VLP (Lin et al., 2022), which uses visual and text information, by $2.3\%p\uparrow$.
@@ 6 Conclusion
Broader impact
In this work, we propose EVEREST and compare it with strong video representation learning baselines on multiple benchmarks. Our goal is to improve training/inference efficiency in video understanding. Our work could have potential societal consequences, none of which we feel must be specifically highlighted here.
@@
The supplementary file is organized as follows: Firstly, we explain the implementation details according to tasks in Appendix A and perform the distance function ablation study for redundancy-robust masking during pre-training in Appendix B. We provide additional experimental results for our EVEREST in Appendix C. Next, we visualize more examples from EVEREST in Appendix D. Finally, we provide a discussion on broader impacts and limitations of our work in Appendix E.
@@ Appendix B Comparing Functions for Computing Token Embedding Distance
Table 10: The ratio of our ReRo Masking at pre-training (Left) and fine-tuning (Right) on UCF101. We highlight the default redundancy-robust masking rate ($\rho_{pre}$) as red texts. We basically pre-train our model 800 epochs, but also report the results with 3,200 pre-training epochs following VideoMAE, denoted by ${}^{\dagger}$. ${}^{\ast}$ and ${}^{\ast\ast}$ denotes the results with 75% and 90% masking, respectively.
Method $\rho_{pre}$ top-1 GFLOPs MAE - - 46.10 ${}^{\dagger}$ ${}^{\ast}$ 57.50 - ${}^{\ast\ast}$ 35.48 Ours 0.5 90.91 23.43 0.4 91.51 21.54 0.3 91.56 19.81 0.2 90.80 18.22 ${}^{\dagger}$ 0.3 93.39 19.81
@@ Appendix D Visualization of Redundancy-Robust Masking and Sampling
Appendix E Broader Impacts and Limitations
Our work has a positive broader societal impact leading to environment-friendly AI research by drastically reducing energy and carbon costs in the training and inference phases. Current methods in the video understanding field consume tremendous monetary and environmental costs, and we believe our work contributes to overcoming this limited feasibility of video-based research fields by enabling single-node training with larger batch sizes using a large backbone, whereas existing VRL methods require immense memory occupancy and suffer from deploying models in the same setup due to out-of-memory. We also aim to enhance the accessibility of those cutting-edge video models by employing practical resources on a single node and pursuing eco-friendly AI research.
On the other hand, while our EVEREST achieves impressive performance improvements along with drastic memory, computation, and training time reduction over multiple benchmark datasets/tasks with different backbone sizes, the improvement in model accuracy for the action recognition task on the K400 dataset becomes relatively small for huge backbones such as ViT-L. We speculate that this is due to their better ability of them to encode spatiotemporal representations into higher-dimensional latent spaces. In addition, our method focuses on capturing short-term motion changes in videos by measuring the temporal disparity between adjacent frames. Although we validated the effectiveness of our method on egocentric videos focusing on temporal semantics or causal interactions, the effect of our method on tasks of long-term episodic memory or the cameraman’s intention/interactions under a long horizon view was not thoroughly analyzed in this study. We leave it to future research to analyze and address these limitations.

### v5

# Removing Redundant Spatiotemporal Tokens

@@ Abstract
(a) Computation (GFLOPs) Comparison
(b) Memory comparison
Figure 1: Efficiency of our EVEREST against VideoMAE on K400 dataset. (a) GFLOPs for pre-training and fine-tuning. The bubble size is proportional to the GFLOPs of the model. (b) Memory consumption using one node equipped with 8$\times$ A100 (80GB). VideoMAE with ViT-B and -L fails to deploy the model due to out-of-memory if the batch sizes are 512 or larger. For a ViT-L backbone with a batch size of 256, our method achieves about 4×4$\times$4 × less memory consumption than VideoMAE. Please see  Tabs. 1 and 5 for detailed results.
@@ 1 Introduction
Massive video data floods the web and media daily with the rapid growth of portable devices equipped with cameras, such as AR glasses, smartphones, UAVs, and robots. However, direct utilization of user-generated video data to solve target task problems is nontrivial, as annotating them is time-consuming and expensive. One potential approach to tackle this problem is to learn generic representations from unlabeled video streams that can transfer to diverse downstream visual tasks. Video Representation Learning (VRL) (Fernando et al., 2017; Piergiovanni et al., 2020; Qian et al., 2021; Pan et al., 2021) methods allow learning spatial and temporal features from input video frames in a self-supervised manner without any human annotations. A caveat to such pre-training for video tasks is that, unlike image-based tasks with static information of objects in instances, video-based tasks involve temporal causality; that is, successive frames are strongly correlated in their semantics.
@@
Recently, Masked Video Autoencoder (MVA) (Tong et al., 2022; Feichtenhofer et al., 2022), which learns to reconstruct randomly masked spatiotemporal regions in video clips, has shown impressive performance on various video-based problems. Yet, critical challenges remain in efficiently exploiting the spatiotemporal information in real-world videos: (1) Tokens (a pair of two temporally successive patches in the same space) in videos are not equally valuable to reconstruct, as the amount of their information depends not only on spatial importance but also on temporal redundancy. (2) Learning representations from videos is infeasible without a huge computing budget. MVA approaches (Feichtenhofer et al., 2022; Wang et al., 2023b, c) that reconstruct the whole video require excessively large amounts of computations, making it impractical without access to a substantial GPU cluster. For example, VideoMAE (Tong et al., 2022) takes about 27 hours to pre-train for 800 epochs with ViT-B using 64 NVIDIA V100 (32GB) GPUs, and ST-MAE (Feichtenhofer et al., 2022) takes about 35.8 hours to pre-train for 800 epochs with ViT-L using 128 NVIDIA A100 (80GB) GPUs.
@@
To overcome such limited feasibility of video representation learning, we propose an Efficient Masked Video AutoEncoder which removes REdundant Spatiotemporal Tokens (EVEREST), that is a highly efficient training algorithm with token selection to alleviate the extremely high resource requirements for VRL. Unlike previous methods (Tong et al., 2022; Feichtenhofer et al., 2022; Wang et al., 2023a) reconstructing all patches, we aim to select a subset of informative visual tokens to learn based on the distance disparity across temporally adjacent tokens in the embedding space. Our redundancy-robust token selection approach successfully detects meaningful changes in the state/object of incoming videos in an online manner, discarding less meaningful tokens containing redundant information or meaningless backgrounds, without resorting to dense motion features from the incoming video, such as HOG or optical flows (Sun et al., 2023; Wang et al., 2023c).
@@
This allows us to significantly reduce computational cost and memory usage while maintaining the quality of a representation model by backpropagating to only a few selected tokens that retain rich spatiotemporal information with minimal redundancy. As shown in Fig. 1, our EVEREST saves computation costs by $\bf 26\sim 45$in pre-training and $\bf 44\sim 48$in fine-tuning across varying ViT scales while achieving competitive performance against strong VideoMAE baselines. We also achieve an impressive amount of memory reduction with all ViT backbones, which becomes more effective with larger models. For example, on a ViT-Large with a batch size of 256, our method achieves about $\bf 4\times$ smaller memory consumption than VideoMAE and enables single-node training with larger batch sizes using a large backbone, whereas existing VRL methods require immense memory occupancy and fail to train in the same setup as they go out-of-memory.
@@ 4.1 Redundancy-robust Mask Generation
$arg\min_{\mathcal{W},\bm{w}}~{}\sum^{N}_{n=1}\sum^{\tau-1}_{i=0}\left\|f_{\mathcal{W}}({(\widetilde{\bm{k}}_{i}^{\prime})^{n}})-{{\bm{m}_{i}^{n}\otimes\bm{v}_{n}}}[2i:2i+1]\right\|_{F}$ (3)
where $\bm{m}^{n}=G\left([J\cdot\rho_{pre}],\tau\right)\in\{0,1\}^{\tau\times{J}}$ and the size $|\bm{m}^{n}|=[J\cdot\tau\cdot\rho_{pre}]$. We omit the positional embedding term for notational simplicity. Whereas earlier MVA works used a fixed masking ratio per video cues for self-supervised training, our approach allows a dynamic masking rate for each frame by design, based on the occupancy of essential tokens. As shown in Fig. 7, each video frame contains a different amount of spatiotemporal information, and our dynamic masking strategy enables the model to focus on learning the valuable representation in a holistic view from incoming videos. We note that our EVEREST can be generalized well on both pre-training and fine-tuning phases with a negligible additional computational cost and training time (Please see Fig. 1). We find that removing unimportant tokens is crucial for both phases, obtaining more meaningful representations and achieving superior performances. (Please see Tab. 10). The overall process of our redundancy-robust token selection is illustrated in Fig. 2.
Figure 3: Information-intensive frame selection. We adaptively select frames based on the ReRo token frequency, which indicates significance compared to frames.
@@
4.2 On-the-fly Information-intensive Frame Selection
Table 1: Performance Comparison on K400. PT and FT mean pre-training and fine-tuning, respectively. We use an input size of $16\times 224^{2}$. Memory usage is measured with 256 batch size when pre-training. is trained from random initialization. VideoMAE and EVEREST are pre-trained for 200 epochs, and we refer (Wightman, 2019) for computing .
@@ 4.2 On-the-fly Information-intensive Frame Selection
MViT (Fan et al., 2021) MViT-S - 32.9 - 76.0
MViT (Fan et al., 2021) MViT-B - 70.5 - 78.4
ViViT FE (Arnab et al., 2021) ViT-L IN-21K 119.0 3980.0 N/A 81.7
K-centered (Park et al., 2022) XViT IN-1K 67.4 425.0 N/A 73.1
K-centered (Park et al., 2022) Mformer IN-1K 67.4 369.5 N/A 74.9
K-centered (Park et al., 2022) TSformer IN-1K 67.4 590.0 N/A 78.0
@@
5 Experiments
Experimental settings
Table 3: Memory usage comparison during the pre-training (one node, A100$\times$8 GPUs) on K400. The memory gaps grow up when increasing batch size and model size. OOM indicates out-of-memory.
@@ 5 Experiments
Table 5: EVEREST-Finetuning with other MVAs on K400. We apply our EVEREST for finetuning the pre-trained models (ViT-B) by SoTA MVAs. We measure the memory usage with the same batch size of 128. While SoTA methods use full tokens during fine-tuning, EVEREST uses only redundancy-robust tokens.
@@
EVEREST has a potential for estimating motion information. Our method is also effective for egocentric videos, e.g., Ego4D, which capture dynamic views from camera motion. EVEREST captures informative objects even with the rapid change of view in first-person videos, as it masks out objects crucial to understanding the camera wearer’s attention while not attending to the background. Specifically, we provide masked input examples with their importance heatmaps through our proposed redundancy-robust masking in Fig. 7. Our masking strategy enables the model to capture the most informative tokens containing static and dynamic moving of objects (e.g., doing wall push up (Fig. 7 (a)) and playing ping pong (Fig. 7 (b)). In addition, it captures informative objects even containing drastic viewpoint changes in videos, as it masks objects crucial for understanding the camera wearers’ attention (Fig. 7 (c)), while not focusing on backgrounds such as walls and floors.
Furthermore, it is worth noting that egocentric videos contain temporally redundant visual information. Fig. 8 reveals that a worker focuses on a grass mower to operate it well, and a person makes cookie dough, where visual scenes include much meaningless visual information, such as empty space on the table. Our method automatically discards these noisy and redundant tokens in an online manner without requiring additional dense information, such as optical flows of incoming videos.
These abundant redundant scenes over a long time in real-world videos market VRL models to waste time and computations by learning on meaningless frames, which also can lead to poor local optima due to bias and catastrophic forgetting. To overcome the limitation, we adopt the redundancy robust sampling rate $\alpha$ to focus on frames with larger motions across frames (See Fig. 4). By constructing the given videos with fluent motion information, our model focuses more on learning the core part of the video. We set the default $\alpha$ to 1.5 and report the effect of $\alpha$ in Appendix C. We quantitatively compare our method with recent SoTA methods using the OSCC task on Ego4D in Tab. 6. We pre-train OSCC without labels and fine-tune it to classify whether the object’s state has changed. With only visual information, we outperform the previous SoTA method, Egocentric VLP (Lin et al., 2022), which uses visual and text information, by $2.3\%p\uparrow$.
@@ 6 Conclusion
Impact Statement
Our work has a positive broader societal impact leading to environment-friendly AI research by drastically reducing energy and carbon costs in the training and inference phases. Current methods in the field of video understanding consume enormous monetary and environmental costs, and we believe that our work contributes to overcoming this limited feasibility of video-based research fields by enabling single-node training with larger batch sizes using a large backbone, whereas existing VRL methods require immense memory occupancy and suffer from deploying models in the same setup due to out-of-memory. We also aim to enhance the accessibility of those cutting-edge video models by employing practical resources on a single node and pursuing eco-friendly AI research.
Acknowledgement
This work was supported by Institute for Information & communications Technology Promotion (IITP) grant funded by the Korea government (MSIP) (No.2019-0-00075 Artificial Intelligence Graduate School Program (KAIST)), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2022-00187238, Development of Large Korean Language Model Technology for Efficient Pre-training), the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (NO. RS-2023-00256259), KAIST-NAVER Hypercreative AI Center, Google Research Grant, and Google Cloud Research Credits program with the award (KX3L-GHDM-4U6F-QLE0).
@@ Acknowledgement
The supplementary file is organized as follows: Firstly, we explain the implementation details according to tasks in Appendix A and perform the distance function ablation study for redundancy-robust masking during pre-training in Appendix B. We provide additional experimental results for our EVEREST in Appendix C. We visualize more examples from EVEREST in Appendix D.
@@ Appendix B Comparing Functions for Computing Token Embedding Distance
Table 10: The ratio of our ReRo Masking at pre-training (Left) and fine-tuning (Right) on UCF101. We highlight the default redundancy-robust masking rate ($\rho_{pre}$) as red texts. We basically pre-train our model 800 epochs, but also report the results with 3,200 pre-training epochs following VideoMAE, denoted by . and ∗∗ denotes the results with 75% and 90% masking, respectively.
Method $\rho_{pre}$ top-1 GFLOPs
MAE - - 46.10
VideoMAE - 91.25 57.50
- 90.80∗∗ 35.48
Ours 0.5 90.91 23.43
0.4 91.51 21.54
0.3 91.56 19.81
0.2 90.80 18.22
Ours 0.3 93.39 19.81